OpenStudy (anonymous):

A manufacturer wants to design an open-top box having a square base and a surface area of 90 square inches. What dimensions will provide a box with maximum volume?

OpenStudy (anonymous):

Please help!!

OpenStudy (anonymous):

Let s and h be the square's edge length and the box height respectively. Then:\[\left\{V\text{=}s^{2 }h,s^{2 }+4s h=90\right\}\]Solve the second equation for h and plug the result back in expression for V. Can you take it from here?

OpenStudy (anonymous):

\[V\text{= }\frac{1}{4} s \left(90-s^2\right)\]The above equation expresses the box volume as a function of s, the square's side length. A Plot is attached.

OpenStudy (anonymous):

thank you very much! @robtobey

OpenStudy (anonymous):

This problem can be solved with one of Mathematica's Controlled Optimization functions, Maximize. No knowledge of the Calculus is necessary. \[\text{Maximize}\left[\left\{s^{2 }h,s^{2 }+4s h=90,s>0,h>0\right\},\{s,h\}\right] \]\[\left\{15 \sqrt{30},\left\{s\to \sqrt{30},h\to \sqrt{\frac{15}{2}}\right\}\right\} \]